OpenStudy (anonymous):

What Is The Quotient Written In scientific notation? 3x10^20 ________ 4x10^5 A. 7.5 x 10^3 B. 0.75 x 10^4 C. 7.2 x 10^14 D. 7.5 x 10^15

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[\huge \frac{3\times 10^{20}}{4\times 10^5}\]

OpenStudy (anonymous):

hi! and yes

OpenStudy (misty1212):

ok first off \[\frac{3}{4}=.75\] right?

OpenStudy (anonymous):

yep

OpenStudy (misty1212):

and \[\frac{10^{20}}{10^5}=10^{20-5}=10^{15}\]

OpenStudy (misty1212):

so one answer (not the final answer though) is \[.75\times 10^{15}\]

OpenStudy (misty1212):

that is not scientific notation because \(.75<1\) so you have to rewrite it as \[7.5\times 10^{14}\]

OpenStudy (anonymous):

thank u!! :)